OpenStudy (foxythepirate0984):

(−1/3) + (−3/5) @MD152727

OpenStudy (md152727):

((-1 × 5) + (-3 × 3)) / (3 × 5) = (-5 + -9) / 15 = -14/15

OpenStudy (foxythepirate0984):

Thank You

OpenStudy (md152727):

Your welcome